OpenStudy (smarty):

Simplify i ^6 • i^8.

OpenStudy (smarty):

Also, these represent Imaginary Numbers. ( The "i"s

myininaya (myininaya):

Same base Since multiplying add exponents \[i^{14}=(i^2)^7\] ---- What is \[i^2=?\]

OpenStudy (smarty):

i*i?

OpenStudy (cwrw238):

hum - i find these tricky by definition i^2 = -1

OpenStudy (cwrw238):

so its (-1)^7

OpenStudy (smarty):

Thats my final answer?

OpenStudy (smarty):

My choices state i -1 -i 1

OpenStudy (cwrw238):

no -1*-1*-1*-1*-1*-1*-1 = 1 * 1 * 1 * -1 = -1

OpenStudy (smarty):

Oh, i have the actually square it by 7... hmk. Thanks~!

OpenStudy (cwrw238):

of you remember that i^2 = -1 you can do these easily also you can remember that i^4 = 1 so foe example i^5 =i^4 * i = 1 * i = i